# Electronic structure and transport in materials with flat bands: 2D materials and quasicrystals

### v1

@@ 3.2.2 Bernal bilayer graphene (BLG)
    with a coefficient $\alpha$ almost independent on the energy $E$. $\alpha$ value is similar to that of MLG [38].
  • ($iii$)
    For energies in the midgap states, $E=E_{MLG}$, the conductivity is very low, and various anomalous behaviors are obtained as in MLG [38]. Indeed, conduction by the midgap states is very specific to the TB model.
@@ 3.2.3 Twisted bilayer graphene (TBG)
The geometry and the very rich physics of magic-angle TBG has been presented in the section 2.3, in particular their quantum diffusion properties for defect-free systems (Sec. 2.3.4). Here, we analyze the diffusion properties in TBG with a random distribution of resonant defects such as adsorbed atoms or molecules, simulated by vacant atoms, like in MLG and BLG (see introduction of Sec. 3.2.2). We focus on the effect of the angle of rotation $\theta$ between the two graphene layers and we compare the cases in which defects are distributed in one layer only (layer 2) to the cases in which defects are distributed in the two layers (layers 1 and 2) [377, 125].
@@
Figure 15: Electronic properties of TBG with defects (vacancies): (a) Total DOS in (12,13) TBG with $\theta=2.66^{\circ}$ bilayer, $N=1873$) for different concentration $c$ of vacancies. (b) Conductivity $\sigma$ versus inelastic mean-free path $L_{i}$ at energy $E=0$ for 3 TBGs with $c=2$. (c) Microscopic conductivity $\sigma_{M}$ in TBG ($\theta=2.66^{\circ}$). Insert: zoom around Dirac energy $E_{D}=0$. (d,e) $\sigma_{M}(E)$ versus $\theta^{2}$ for $E=0$ and $E=0.1$ eV, respectively. In panels (a,c,d,e): (full line) with vacancies in layer 2 and (dashed line) with vacancies in both layers. SK-TB model is the one which results in an intraband monolayer velocity $V_{MLG}=0.79\cdot 10^{6}$ ms-1 (Sec. 2.3.2). TBGs are rigid, i.e., without atomic relaxation. For the numerical method see A.3. $G_{0}={2e^{2}}/{h}$. From [125].
@@
The effect of static defects (vacancies) on total density of states (DOS) and microscopic conductivity $\sigma_{M}$ (i.e., room-temperature conductivity, Sec. 3.1) are shown in Fig. 15. As expected, when disorder increases, i.e., the concentration $c$ of vacancies increases, the two van Hove singularity are gradually destroyed (Fig. 15(a)).
@@
For vacancy distributions in the two layers $\sigma_{M}(E)$ is almost independent of $\theta$ value (see dashed lines in Fig. 15(d,e)). When vacancy concentration $c$ is large, the behavior is similar to that of MLG and $\sigma_{M}\approx 2\sigma_{M,{MLG}}$, where $\sigma_{M,{MLG}}$ is the conductivity for MLG [38], $\sigma_{M,{MLG}}\approx 0.6$ G0. $\sigma_{M,{MLG}}$ reaches the well known universal minimum of the conductivity so-called conductivity “plateau” –independent of defect concentration– at energies around $E_{D}$ [239]. For smaller concentration $c$, $\sigma_{M}$ increases when $c$ increases. These two regimes are similar to the one found in AB Bernal bilayer graphene [358] (see Sec. 3.2.2). Roughly speaking, for large value of $c$ , the elastic mean free path $L_{e}$ in MLG is smaller than the average travel distance $l_{1}\approx 1.5$ nm within a layer, between two interlayer hoppings of the charge carriers, and thus carriers behave as they would in MLG. Whereas for small $c$ values, $L_{e}>l_{1}$ and thus interlayer hoppings are involved in the diffusive regime and BLG conductivity properties are different that MLG ones.
@@
For vacancies distribution in layer 2 only, and large rotated angle (see solid line in Fig. 15(d,e)), the conductivity is larger than in the previous case. Indeed for large $\theta$, typically $\theta>10^{\circ}$, eigenstates are located mainly in one layer (“decoupled” layers) [90, 29] and thus conductivity of the bilayer is the sum of the conductivity of two almost independent layers,
@@
corresponding to the conductivity of layers 1 and 2, respectively. The conductivity of a layer with defects is close to the MLG conductivity $\sigma_{M,2}\approx\sigma_{M,{MLG}}$ and the conductivity of layer without defects $\sigma_{M,1}$ is affected by the presence of defects in layer 2. With increasing $\theta$, the eigenstates are more and more located on one layer, thus layers are more and more decoupled, and the $\sigma_{M,1}$ increases as layer 1 becomes more and more like a pristine MLG. Consequently the conductivity of the TBG increases when $\theta$ increases. In theses cases numerical results (Fig. 15(d,e)) show that $\sigma_{M}$ increases as $\theta^{2}$. This behavior has been clarified by an continuous analytic model [244, 125], showing that for intermediate angles,
@@
where $\theta_{0}\approx 1.7^{\circ}$ (Eq. (24)), and $n_{1}$, $n_{2}$ are the DOS in layer 1 and 2 that are different since the midgap state peak is not seen in defect-free layer [125].
For small angles, eigenstates are located almost equally on both layers for all energies around Dirac energy [29] (Sec. 2.3.4); therefore they are affected in a similar way by the two kinds of vacancy distributions: ($i$) in the two layers and ($ii$) in layer 2. Conductivity is thus very similar in the two cases.
In the localization regime (i.e., the low-temperature limit, $L_{i}\gg L_{e}$), the multiple scattering effects may reduced the conductivity with respect to microscopic conductivity $\sigma_{M}$. The inelastic length $L_{i}$ thus satisfies $L_{i}\gg L_{e}$ and $L_{i}\gg l_{1}$. To evaluate this effect we compute the conductivity $\sigma$ versus $L_{i}$ at every energy $E$ [125] for the two vacancy distribution cases: ($i$) and ($ii$). As expected in disordered 2D systems [189], for large $L_{i}$, $\sigma(L_{i})$ follows a linear variation with the logarithm of $L_{i}$ (Eq. 34), as in the case of MLG [38, 371] and BLG (Sec. 3.2.2). Similary to $\sigma_{M}$, for intermediated angle TBG, the localization length $\xi$ is almost independent of $\theta$ when the defects are located in both layers, but $\xi$ increases sharply when defects are located in a single layer. Despite some studies, e.g. [124], further studies are needed to understand the effect of resonant defects on the magic-angle TBG properties, in particular to find out how it can destroy the moiré flat bands.
@@ 3.3 Selective distribution of defects in bipartite systems
As already mentioned, the remarkable electronic properties of MLG and BLG make then a good candidate for many applications, but these applications are severely limited by the absence of a gap. Hence, the band-gap opening and the control of graphene bilayer become essential for the applications in various electronic devices. One way to create a gap in MLG is the selective functionalization. It has been done, for example, with hydrogen adsorption on a moiré pattern of Graphene-Ir(111) [87]. BLG offers more possibilities for obtaining a gap. One of its advantages is the control of its gap by applying an external gate voltage [378, 379, 208], which opens the way to multiple applications for nanodevices [380, 381, 382]. On the other hand, the BLG electrical conductivity can be performed with using the influence of substrate [383], vacancies, adatoms or admolecules adsorbed on the surface of BLG [384, 385, 349, 358, 354, 386, 349, 387, 350, 388, 160]. For instance, single- and double-sided fluorination has been shown to affect strongly conductivity, exhibiting insulating and conducting behavior, respectively [350]. This suggests that the specific functionalization of only one of the two sublattice should be an also important parameter for controlling a real gap or a mobility gap [349, 354, 388, 160].
In this section, we study the effect of a selective functionalization in a bipartite lattice. “Selective functionnalization” means that defects (vacancies or adsorbates) are randomly distributed in a specific sublattice of the system. MLG and BLG are bipartite lattices, i.e., their lattice is composed in two sublattices, $\alpha$ and $\beta$, such as all first neighbors of an $\alpha$ atom is a $\beta$ atom and conversely. In Sec. 3.3.1, we focus on density of states (DOS) when defects are located only on $\alpha$ sublattice or $\beta$ sublattice. Indeed in this case an effective Hamiltonian can be written to simply understand the effect of the defects. In Sec. 3.3.2, we present a detailed numerical study of BLG transport properties when one or two types of adsorption site are present simultaneously. In Sec. 3.3.3, we focus on the very particular conductivity of midgap states at finite temperature.
@@ 3.3.1 Effective Hamiltonian for defects in a single sublattice of a bipartite lattice
Considering in this section a bipartite lattice and the first-neighbor TB Hamiltonian $\hat{H}$ (Eq. 33), $\hat{H}$ couples only $\alpha$ states ($\beta$ states) with $\beta$ states ($\alpha$ states). Quite generally an eigenstate $\varphi$ with energy $E$ of $\hat{H}$ (with or without vacancies), can be written as $\varphi=\varphi_{\alpha}+\varphi_{\beta}$, with states $\varphi_{\alpha}$ ($\varphi_{\beta}$) belonging to the sublattice $\alpha$ ($\beta$). It is easy to show that $\varphi_{\alpha}$ and $\varphi_{\beta}$ are eigenstates of the effective Hamiltonian,
@@
For non-zero energy the DOS $\tilde{n}(\tilde{E})$ of $\hat{\tilde{H}}$ is related to the DOS $n_{\alpha}(E)$ and $n_{\beta}(E)$ on sub-parts $\alpha$ and $\beta$ by,
$n_{\alpha}(E)=n_{\beta}(E)=2\,|E|\,\tilde{n}(\tilde{E}).$ (39)
These relations show that the spectrum of a bipartite lattice with first-neighbor TB coupling is necessarily symmetrical with respect to the on-site energy of the orbitals (here $E_{D}=0$). This symmetry is lifted if we introduce TB couplings beyond the nearest neighbours, but the results we present in the following still give good qualitative behavior, as we have verified in the case of the BLG [354].
We now analyze the average local density of states $\tilde{n}_{i}(\tilde{E})$ (LDOS) of $\hat{\tilde{H}}$ in presence of vacancies located in a single sublattice, where $i$ is the index of a sublattice. Vacancies in a single sublattice ($\alpha$ or $\beta$) creates midgap states at $E=0$ that are not coupled together [351, 81]. Thoses midgap states at $E=0$ are not drawn in DOS shown in this section.
@@
Figure 16: Monolayer graphene (MLG) with vacancies in the A sublattice: (a) A sublattice in MLG (without vacancy) with the sketch of $\hat{\tilde{H}}=\hat{H}^{2}$ hopping parameters, where $\hat{H}$ it the first-neighbor TB Hamiltonian (hopping term $t_{0}=2.7$ eV). The A sublattice is a triangular lattice. (b) Total density of states DOS $n$ of $\hat{H}$ (black line is without vacancies $c=0$). The midgap state at $E=0$ (located on B sublattice) are not shown. Note that the average local DOS on sublattice A and B are the same for $E\neq 0$ (Eq. (39)), therefore per atom: $n=n_{A}=n_{B}({\rm without~{}peaks~{}of~{}midgap~{}states~{}at~{}}E=0)$. (c) Average local DOS on triangular sublattices A, $\tilde{n}_{A}(\tilde{E}=E^{2})$, of the Hamiltonian $\hat{\tilde{H}}=\hat{H}^{2}$; (c’) is a zoom of (c) at low energies for the same concentration of vacancies as in panel (b). (d) Minimum value of $\hat{H}^{2}$ spectrum, $\tilde{E}_{m}=E^{2}_{m}$, versus $c$ (the line is a guide for the eyes). $c$ is the concentration of vacancies with respect to the total number of atoms in MLG. Numerical method see A.3. From [354, 376].
@@
For MLG [354], $\alpha$ states are $p_{z}$ orbitals of A atoms and $\beta$ states are $p_{z}$ orbitals of B atoms. The effective Hamiltonian ${\hat{\tilde{H}}}$ is a Hamiltonian of a triangular lattice of $\alpha={\rm A}$ sublattice,
@@
where $t_{0}=2.7$ eV is the first neighbor hopping term. The hopping terms $\tilde{t}_{0}$ are sketched Fig. 16(a). The middle of $\hat{H}$ band ($E=0$) corresponds to lowest energy of $\hat{\tilde{H}}$ band ($\tilde{E}=0$). $n(E)$ calculated numerically from $\hat{H}$ is shown in Fig. 16(b), and $\tilde{n}(\tilde{E})$ calculated from $\hat{\tilde{H}}$ is shown in Fig. 16(c). These results verify equation (39).
The effect of vacancies in sublattice ${\rm A}$ on the DOS away from zero energy can be understood [354] by considering $\hat{\tilde{H}}_{{\rm A}}$ (Eq. 40), but with functionalized ${\rm A}$ sites which are simply deleted. Without vacancies the coordination $\eta$ of each atoms of ${\rm A}$ sublattice is 6. With a small concentration $c$ of vacancies in the ${\rm A}$ sublattice, the average coordination is,
@@
The center of the ${\rm A}$ band is fixed by on-site energies $\tilde{\epsilon}_{{\rm A}}$ and it is not affected by vacancies; but the width of the band will decrease when $\eta$ decreases (i.e., when $c$ increases) (see Figs. 16(c,c’)). As expected from this simple tight-binding argument, the minimum values, $\tilde{E}_{m}=E^{2}_{m}$, of the spectrum of $\hat{\tilde{H}}$, found numerically (Fig. 16(d)), is almost proportional to the average coordination number $\eta$ (average number of A-A (B-B) nearest neighbors of A (B) sublattice of the bipartite lattice). Consequently the average ${\rm A}$ DOS, $\tilde{n}_{{\rm A}}$, has a gap induced by vacancies for $0\leq\tilde{E}\leq\tilde{E}_{m}$. This means that the DOS in the $\alpha={\rm A}$ and $\beta={\rm B}$ sublattices of MLG also presents a gap for $-E_{m}\leq E\leq E_{m}$, with ${E}_{m}=\sqrt{\tilde{E}_{m}}$ (Fig. 16(b)).
@@
Figure 17: BLG with vacancies in the $\alpha$ sublattice: (a) $\alpha=\{{\rm A}_{1};{\rm B}_{2}\}$ sublattice in BLG (without vacancy) with $\hat{H}^{2}$ hopping parameters. (b) Average LDOS of $\hat{\tilde{H}}=\hat{H}^{2}$ of pristine BLG on sublattice A1, $\tilde{n}_{{\rm A}_{1}}(\tilde{E})$, and on sublattice B2, $\tilde{n}_{{\rm B}_{2}}(\tilde{E})$. (dashed line) show LDOS without A1-B2 hopping, $\tilde{t}_{{\rm A}_{1}{\rm B}_{2}}=0$. Average LDOS on (c) atoms A1, $\tilde{n}_{{\rm A}_{1}}(\tilde{E})$, and (e) atoms B2, $\tilde{n}_{{\rm B}_{2}}(\tilde{E})$, with vacancies on A1 sublattice. Average LDOS on (d) atoms A1, $\tilde{n}_{{\rm A}_{1}}(\tilde{E})$, and (f) atoms B2, $\tilde{n}_{{\rm B}_{2}}(\tilde{E})$, with vacancies on B2 sublattice. $c$ is the concentration of vacancies with respect to the total number of atoms in BLG. The corresponding DOS of $\hat{H}$, $n(E)$, and the microscopic conductivity are shown Fig. 18(a,b). Numerical method see A.3. From [376].
@@
where $\langle i,j\rangle$ represents the corresponding nearest-neighbor pairs of occupied sites. $\hat{\tilde{H}}_{{\rm A}_{1}}$ and $\hat{\tilde{H}}_{{\rm B}_{2}}$ are Hamiltonian of A1 and B2 states respectively; they are coupled by $\hat{\tilde{H}}_{{\rm A}_{1}{\rm B}_{2}}$. The coupling parameters of these Hamiltonians are deduced from coupling parameters in BLG first-neighbour TB Hamiltonian (intralayer first neighbor coupling $t_{0}=2.7$ eV and the interlayer first neighbor coupling $t_{1}=0.34$ eV, see Sec. 3.2). For the triangular sublattice A1, the on-site energies are
@@
The local density of states without vacancies $\tilde{n}_{{\rm A}_{1}}(\tilde{E})$ and $\tilde{n}_{{\rm B}_{2}}(\tilde{E})$ on an atom A1 and on an atom B2, respectively, are shown in Fig. 17(b). If the A1-B2 coupling is turned off, $\tilde{t}_{1}=0$, while maintaining $t_{1}\neq 0$ in Eq. (47), (see dashed lines in Fig. 17(b)), $\tilde{n}_{{\rm A}_{1}}(\tilde{E})$ and $\tilde{n}_{{\rm B}_{2}}(\tilde{E})$ correspond to two independent bands; at low energies they are constant for $\tilde{E}\geq\tilde{\epsilon}_{{\rm A}_{1}}-3\tilde{t}_{{\rm A}_{1}{\rm A}_{1}}=t_{1}^{2}=0.1156$ eV2 and $\tilde{E}\geq\tilde{\epsilon}_{{\rm B}_{2}}-3\tilde{t}_{B_{\rm 2}{\rm B}_{2}}=0$, respectively. In that case, from Eq. (39), $n_{MLG}(E)=2{|E|}\tilde{n}_{{\rm B}_{2}}(E^{2})$ is the DOS of monolayer graphene (MLG).
We now consider the case with selective resonant defects (vacancies). Similarly to MLG case, vacancies in A1 sublattice (or B2 sublattice) create a pseudo-gap at the lowest energies of A1 DOS (B2 DOS) (Fig. 17(c,f)). Note that this pseudo-gap is a real gap for $t_{1}=0$ (MLG case). In BLG, because of this pseudo-gap the low-energy DOS of the clean sublattice ($\beta$) recovers its value in the absence of the effective hopping $\tilde{t}_{1}=\tilde{t}_{{\rm A}_{1}{\rm B}_{2}}$. Therefore the eigenstates of the clean sublattice are simply obtained by suppressing the defected sublattice. According to the basic relations in bipartite Hamiltonian explained above, the component of the eigenstate in sub-part $\beta=\{{\rm A}_{2},{\rm B}_{1}\}$ is also obtained by suppressing the contribution of the defected sublattice. Therefore our main conclusion is that the electronic structure of the BLG in the low energy region where the pseudo-gap occurs in the defected sublattice is simply given by deleting the defected sublattice. We can consider two extreme cases [354]:
@@
These results are also consistent with the those obtained from numerical calculation performed with a Hamiltonian including hopping terms beyond nearest neighbors [354].

### v2

@@ 3.2.2 Bernal bilayer graphene (BLG)
    with a coefficient $\alpha$ almost independent of the energy $E$. The value of $\alpha$ is similar to that of MLG Trambly13 .
  • (iiiiii)
    For energies in the midgap states, the conductivity is very low, and various anomalous behaviors are obtained as in MLG Trambly13 (see also Sec. 3.3.3). Indeed, conduction by the midgap states is very specific to the TB model.
@@ 3.2.3 Twisted bilayer graphene (TBG)
The geometry and the very rich physics of magic-angle TBG has been presented in the section 2.3, in particular their quantum diffusion properties for defect-free systems (Sec. 2.3.4). Here, we analyze the diffusion properties in TBG with a random distribution of resonant defects such as adsorbed atoms or molecules, simulated by vacant atoms, like in MLG and BLG (see introduction of Sec. 3.2.2). We focus on the effect of the angle of rotation $\theta$ between the two graphene layers and we compare the case in which defects are distributed in one layer only (layer 2) to the case in which defects are distributed in the two layers (layers 1 and 2) Toai_these , Omid20 .
@@
Figure 15: Electronic properties of TBG with defects (vacancies): (a) Total DOS in (12,13) TBG ($\theta=2.66^{\circ}$, $N=1873$) for different concentration $c$ of vacancies. (b) Conductivity $\sigma$ versus inelastic mean free path $L_{i}$ at energy $E=0$ for 3 TBGs with $c=2$ %. (c) Microscopic conductivity $\sigma_{M}$ in TBG ($\theta=2.66^{\circ}$). Insert: zoom around the Dirac energy $E_{D}=0$. (d),(e) $\sigma_{M}(E)$ versus $\theta^{2}$ for $E=0$ and $E=0.1$ eV, respectively. In all panels: (full lines) with vacancies in layer 2 and (dashed lines) with vacancies in both layers. The SK-TB model is the one which results in an intraband monolayer velocity $V_{MLG}=0.79\cdot 10^{6}$ ms-1 (Sec. 2.3.2). TBGs are rigid, i.e., without atomic relaxation. For the numerical method see A.3. $G_{0}={2e^{2}}/{h}$. From Omid20 .
@@
The effect of static defects (vacancies) on total density of states (DOS) and microscopic conductivity $\sigma_{M}$ (i.e., room-temperature conductivity, Sec. 3.1) are shown in Fig. 15. As expected, when disorder increases, i.e., the concentration $c$ of vacancies increases, the two van Hove singularities are gradually destroyed (Fig. 15(a)).
@@
For vacancy distributions in the two layers $\sigma_{M}(E)$ is almost independent of the value of $\theta$ (see dashed lines in Fig. 15(d,e)). When the vacancy concentration $c$ is large, the behavior is similar to that of MLG and $\sigma_{M}\approx 2\sigma_{M,{MLG}}$, where $\sigma_{M,{MLG}}$ is the conductivity for MLG Trambly13 , $\sigma_{M,{MLG}}\approx 0.6$ G0. $\sigma_{M,{MLG}}$ reaches the well-known universal minimum of the conductivity so-called conductivity “plateau” –independent of defect concentration– at energies around $E_{D}$ Castro09_RevModPhys . For smaller concentration $c$, $\sigma_{M}$ increases when $c$ increases. These two regimes are similar to the one found in AB Bernal bilayer graphene Missaoui17 (see Sec. 3.2.2). Roughly speaking, for large values of $c$ , the elastic mean free path $L_{e}$ in MLG is smaller than the average traveling distance $l_{1}\approx 1.5$ nm within a layer, between two interlayer hoppings of the charge carriers, and thus carriers behave as they would in MLG. Whereas for small $c$ values, $L_{e}>l_{1}$ and thus interlayer hoppings are involved in the diffusive regime and BLG conductivity properties are different from MLG ones.
@@
For vacancies distribution in layer 2 only, and large rotation angle (see solid lines in Fig. 15(d,e)), the conductivity is larger than in the previous case. Indeed for large $\theta$, typically $\theta>10^{\circ}$, eigenstates are located mainly in one layer (“decoupled” layers) Trambly10 , Trambly16 and thus conductivity of the bilayer is the sum of the conductivities of two almost independent layers,
@@
corresponding to the conductivity of layers 1 and 2, respectively. The conductivity of a layer with defects is close to the MLG conductivity $\sigma_{M,2}\approx\sigma_{M,{MLG}}$ and the conductivity of the layer without defects $\sigma_{M,1}$ is affected by the presence of defects in layer 2. With increasing $\theta$, the eigenstates are more and more located on one layer, thus layers are more and more decoupled, and $\sigma_{M,1}$ increases as layer 1 becomes more and more like a pristine MLG. Consequently the conductivity of the TBG increases when $\theta$ increases. In these cases numerical results (Fig. 15(d,e)) show that $\sigma_{M}$ increases as $\theta^{2}$. This behavior has been clarified by a continuous analytic model Omid_these , Omid20 , showing that for intermediate angles,
@@
where $\theta_{0}\approx 1.7^{\circ}$ (Eq. (24)), and $n_{1}$, $n_{2}$ are the DOS in layer 1 and 2 that are different since the midgap state peak is not seen in a defect-free layer Omid20 .
For small angles, eigenstates are located almost equally on both layers for all energies around the Dirac energy Trambly16 (Sec. 2.3.4); therefore they are affected in a similar way by the two kinds of vacancy distributions: ($i$) in the two layers and ($ii$) in layer 2. Conductivity is thus very similar in the two cases.
In the localization regime (i.e., the low-temperature limit, $L_{i}\gg L_{e}$), the multiple scattering effects may reduce the conductivity with respect to the microscopic conductivity $\sigma_{M}$. The inelastic length $L_{i}$ thus satisfies $L_{i}\gg L_{e}$ and $L_{i}\gg l_{1}$. To evaluate this effect we compute the conductivity $\sigma$ versus $L_{i}$ at every energy $E$ Omid20 for the two vacancy distribution cases: ($i$) and ($ii$). As expected in disordered 2D systems Lee85 , for large $L_{i}$, $\sigma(L_{i})$ follows a linear variation with the logarithm of $L_{i}$ (Eq. (34)), as in the case of MLG Trambly13 , Trambly11 and BLG (Sec. 3.2.2). Similarly to $\sigma_{M}$, for intermediate-angle TBG, the localization length $\xi$ is almost independent of $\theta$ when the defects are located in both layers, but $\xi$ increases sharply when defects are located in a single layer. Despite some studies, e.g., Andelkovic18 , further studies are needed to understand the effect of resonant defects on the magic-angle TBG properties, in particular to find out how it can destroy the moiré flat bands.
@@ 3.3 Selective distribution of defects in bipartite systems
As already mentioned, the remarkable electronic properties of MLG and BLG make them good candidates for many applications, but these applications are severely limited by the absence of a gap. Hence, the band-gap opening and the control of bilayer graphene become essential for the applications in various electronic devices. One way to create a gap in MLG is the selective functionalization. It has been done, for example, with hydrogen adsorption on a moiré pattern of graphene-Ir(111) Jorgensen16 . BLG offers more possibilities for obtaining a gap. One of its advantages is the control of its gap by applying an external gate voltage Castro07 , McCann06 , McCann13 , which opens the way to multiple applications for nanodevices Zhang09 , Overweg18 , Kurzmann19 . On the other hand, the BLG electrical conductivity can be tuned with using the influence of substrate Zhou07 , vacancies, adatoms or admolecules adsorbed on the surface of BLG Leenaerts09 , Mapasha12 , VanTuan16 , Missaoui17 , Missaoui18 , Katoch18 , VanTuan16 , Pinto20 , Son20 , Joucken21 , Jouda21 . For instance, single- and double-sided fluorination has been shown to affect conductivity strongly, exhibiting insulating and conducting behavior, respectively Son20 . This suggests that the specific functionalization of only one of the two sublattices should be also an important parameter for controlling a real gap or a mobility gap VanTuan16 , Missaoui18 , Joucken21 , Jouda21 .
In this section, we study the effect of a selective functionalization in a bipartite lattice. “Selective functionnalization” means that defects (vacancies or adsorbates) are randomly distributed in a specific sublattice of the system. MLG and BLG are bipartite lattices, i.e., their lattice is composed of two sublattices, $\alpha$ and $\beta$, such that any first neighbor of an $\alpha$ atom is a $\beta$ atom and conversely. In Sec. 3.3.1, we focus on the density of states (DOS) when defects are located only on sublattice $\alpha$ or sublattice $\beta$ . In this case an effective Hamiltonian can be written to simply understand the effect of the defects. In Sec. 3.3.2, we present a detailed numerical study of BLG transport properties when one or two types of adsorption site are present simultaneously. In Sec. 3.3.3, we focus on the very particular conductivity of midgap states at finite temperature.
@@ 3.3.1 Effective Hamiltonian for defects in a single sublattice of a bipartite lattice
Considering in this section a bipartite lattice and the first-neighbor TB Hamiltonian $\hat{H}$ (Eq. (33)), $\hat{H}$ couples only $\alpha$ states ($\beta$ states) with $\beta$ states ($\alpha$ states). Quite generally an eigenstate $\varphi$ with energy $E$ of $\hat{H}$ (with or without vacancies), can be written as $\varphi=\varphi_{\alpha}+\varphi_{\beta}$, with states $\varphi_{\alpha}$ ($\varphi_{\beta}$) belonging to the sublattice $\alpha$ ($\beta$). It is easy to show that $\varphi_{\alpha}$ and $\varphi_{\beta}$ are eigenstates of the effective Hamiltonian,
@@
[…]
These relations show that the spectrum of a bipartite lattice with first-neighbor TB coupling is necessarily symmetrical with respect to the on-site energy of the orbitals (here $E_{D}=0$). This symmetry is lifted if we introduce TB couplings beyond the nearest neighbors, but the results we present in the following still give good qualitative behavior, as we have verified in the case of the BLG Missaoui18 .
[…]
@@
Figure 16: Monolayer graphene (MLG) with vacancies in the A sublattice: (a) A sublattice in MLG (without vacancy) with the sketch of $\hat{\tilde{H}}=\hat{H}^{2}$ hopping parameters, where $\hat{H}$ it the first-neighbor TB Hamiltonian (hopping term $t_{0}=2.7$ eV). The A sublattice is a triangular lattice. (b) Total density of states DOS $n$ of $\hat{H}$ (the black line is without vacancies $c=0$). The midgap states at $E=0$ (located on the B sublattice) are not shown. Note that the average local DOS on sublattices A and B are the same for $E\neq 0$ (Eq. (39)), therefore per atom: $n=n_{A}=n_{B}({\rm without~peaks~of~midgap~states~at~}E=0)$. (c) Average local DOS on triangular sublattices A, $\tilde{n}_{A}(\tilde{E}=E^{2})$, of the Hamiltonian $\hat{\tilde{H}}=\hat{H}^{2}$; (c’) is a zoom of (c) at low energies for the same concentration of vacancies as in panel (b). (d) Minimum value of the spectrum of $\hat{H}^{2}$, $\tilde{E}_{m}=E^{2}_{m}$, versus $c$ (the line is a guide for the eyes). $c$ is the concentration of vacancies with respect to the total number of atoms in MLG. For the numerical method see A.3. From Missaoui18 , Missaoui_these .
@@
For MLG Missaoui18 , $\alpha$ states are $p_{z}$ orbitals of A atoms and $\beta$ states are $p_{z}$ orbitals of B atoms. The effective Hamiltonian $^{\hat{\tilde{H}}}$ is a Hamiltonian of a triangular lattice of the $\alpha={\rm A}$ sublattice,
@@
[…]
The effect of vacancies in sublattice $\rm A$} on the DOS away from zero energy can be understood Missaoui18 by considering $\hat{\tilde{H}}_{{\rm A}}$ (Eq. (40)), but with functionalized $\rm A$} sites which are simply deleted. Without vacancies the coordination $\eta$ of each atom of the $\rm A$} sublattice is 6. With a small concentration $c$ of vacancies in the $\rm A$} sublattice, the average coordination is,
@@
The center of the $\rm A$} band is fixed by on-site energies $\tilde{\epsilon}_{{\rm A}}$ and it is not affected by vacancies; but the width of the band will decrease when $\eta$ decreases (i.e., when $c$ increases) (see Figs. 16(c,c’)). As expected from this simple tight-binding argument, the minimum values, $\tilde{E}_{m}=E^{2}_{m}$, of the spectrum of $\hat{\tilde{H}}$, found numerically (Fig. 16(d)), is almost proportional to the average coordination number $\eta$ (average number of A-A nearest neighbors of the A sublattice of the bipartite lattice). Consequently the average $\rm A$} DOS, $\tilde{n}_{{\rm A}}$, has a gap induced by vacancies for $0\leq\tilde{E}\leq\tilde{E}_{m}$. This means that the DOS in the $\alpha={\rm A}$ and $\beta={\rm B}$ sublattices of MLG also presents a gap for $E_{m}\leq E\leq E_{m}$, with ${E}_{m}=\sqrt{\tilde{E}_{m}}$ (Fig. 16(b)).
@@
[…]
@@
where $\langle i,j\rangle$ represents the corresponding nearest-neighbor pairs of occupied sites. $\hat{\tilde{H}}_{{\rm A}_{1}}$ and $\hat{\tilde{H}}_{{\rm B}_{2}}$ are the Hamiltonians of A1 and B2 states respectively; they are coupled by $\hat{\tilde{H}}_{{\rm A}_{1}{\rm B}_{2}}$. The coupling parameters of these Hamiltonians are deduced from coupling parameters in the BLG first-neighbor TB Hamiltonian (intralayer first-neighbor coupling $t_{0}=2.7$ eV and the interlayer first-neighbor coupling $t_{1}=0.34$ eV, see Sec. 3.2). For the triangular sublattice A1, the on-site energies are
@@
[…]
We now consider the case with selective resonant defects (vacancies). Similarly to the MLG case, vacancies in the A1 sublattice (or the B2 sublattice) create a pseudogap at the lowest energies of A1 DOS (B2 DOS) (Fig. 17(c,f)). Note that this pseudogap is a real gap for $t_{1}=0$ (MLG case). In BLG, thanks to this pseudogap the low-energy DOS of the clean sublattice ($\beta$) recovers its value in the absence of the effective hopping $\tilde{t}_{1}=\tilde{t}_{{\rm A}_{1}{\rm B}_{2}}$. Therefore the eigenstates of the clean sublattice are simply obtained by suppressing the defected sublattice. According to the basic relations in a bipartite Hamiltonian explained above, the component of the eigenstate in sub-part $\beta=\{{\rm A}_{2},{\rm B}_{1}\}$ is also obtained by suppressing the contribution of the defected sublattice. Therefore our main conclusion is that the electronic structure of BLG in the low-energy region where the pseudogap occurs in the defected sublattice is simply given by deleting the defected sublattice. We can consider two extreme cases Missaoui18 :
@@
These results are also consistent with the those obtained from numerical calculation performed with a Hamiltonian including hopping terms beyond nearest neighbors Missaoui18 .